What is 35% of £86 and can u show me the workin out for the answer please?
Answer by askmemath(368)

The word "cent" is derived from French meaning 100
So "per cent" means per 100
Now you want 35% of 86
35/100 of 86 = 35/100 x 86 = 3010/100 = 30.1